### v1

# arXiv:2602.23381v1[math.GN] 03 Feb 2026

Abstract
We develop a framework for shallow and deep neural networks whose inputs range over a general topological space. The model is built from a prescribed family of continuous feature maps and a fixed scalar activation function, and it reduces to multilayer feedforward networks in the Euclidean case. We focus on the universal approximation property and establish general conditions under which such networks are dense in spaces of continuous vector-valued functions on arbitrary and locally convex topological spaces. In the absence of width constraints, we obtain universality results that extend classical approximation theorems to non-Euclidean settings. A central focus of the paper is the deep narrow framework, in which the width of each hidden layer is uniformly bounded while the depth is allowed to grow. We identify conditions under which such width-constrained deep networks retain universal approximation power. As a concrete example, we employ Ostrand’s extension of the Kolmogorov superposition theorem to derive an explicit universality result for products of compact metric spaces, with width bounds expressed in terms of topological dimension.
Keywords: Universal approximation property; shallow neural networks; deep neural networks; deep narrow neural networks; topological spaces; locally convex spaces; topological embeddings; topological dimension; Kolmogorov–Ostrand superposition theorem.
2020 MSC: 68T07, 41A46, 26B40, 46E10, 54F45
1 Introduction
Neural networks are a central tool in contemporary machine learning and artificial intelligence. Among the many architectures in use, multilayer feedforward neural networks remain a basic and influential model, both in practice and in theory. A broad overview of the approximation-theoretic properties of feedforward neural networks can be found, for example, in [28, 11], while modern deep learning practice is surveyed in [7].
A multilayer feedforward neural network consists of an input layer, an output layer, and one or more hidden layers arranged sequentially. Information propagates forward through the network: each neuron forms an affine combination of the outputs of the previous layer and then applies a fixed nonlinear activation function. This repeated affine–nonlinear processing enables the network to represent highly complex nonlinear mappings.
The approximation theory of neural networks is rooted in the universal approximation property, also called the density property. In the classical formulation, for suitable activation functions $\sigma$, the linear span of functions of the form $\sigma(\mathbf{w}\cdot\mathbf{x}-\theta)$ is dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$. A fundamental result in this direction states that, when the activation function $\sigma$ is assumed to be continuous, the universal approximation property holds if and only if $\sigma$ is not a polynomial (see [28, 23]).
Note that the inner product $\mathbf{w}\cdot\mathbf{x}$ appearing in classical neural networks is a linear continuous functional on $\mathbb{R}^{d}$. Conversely, by the Riesz representation theorem, every linear continuous functional on $\mathbb{R}^{d}$ can be written in this form (see [30, Theorem 13.32]). Let $\mathcal{L}(\mathbb{R}^{d})$ denote the space of all linear continuous functionals on $\mathbb{R}^{d}$. From this viewpoint, a single hidden layer network applies a collection of functions from $\mathcal{L}(\mathbb{R}^{d})$ to the input, adds scalar biases, and then applies a nonlinearity. In particular, the classical universal approximation property can be rephrased as the density of the function class
$\mathcal{M}(\sigma)=\operatorname{span}\bigl\{\sigma(f(x)-\theta):f\in\mathcal{L}(\mathbb{R}^{d}),\ \theta\in\mathbb{R}\bigr\}$
in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$.
From Euclidean inputs to general topological inputs. The preceding reformulation highlights that, in standard feedforward neural networks, the only structural use of the Euclidean input space $\mathbb{R}^{d}$ lies in the availability of a rich family of scalar-valued continuous functions, namely, linear functionals, applied to the input. This observation suggests a natural extension beyond Euclidean spaces.
Let $X$ be an arbitrary topological space. Instead of linear functionals, we fix a family $\mathcal{A}(X)\subset C(X)$ of continuous real-valued functions on $X$, which play the role of admissible scalar feature maps available to the network. In the classical case $X=\mathbb{R}^{d}$, the choice $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$ recovers exactly the standard architecture, since each feature map $f\in\mathcal{A}(X)$ corresponds to an inner product $\mathbf{w}\cdot\mathbf{x}$.
Definition 1.1 (Basic family).
A set $\mathcal{A}(X)\subset C(X)$ is called a basic family (for feedforward neural networks on $X$) if its elements serve as the admissible scalar feature maps applied directly to the input in the first affine transformation of the network.
We now define shallow topological feedforward neural networks in a form that directly mirrors the standard affine–activation–affine structure of classical single hidden layer networks.
Definition 1.2 (Single hidden layer TFNN with vector output).
Fix a topological space $X$, a basic family $\mathcal{A}(X)\subset C(X)$, an activation function $\sigma:\mathbb{R}\to\mathbb{R}$, and an integer $m\geq 1$. A topological feedforward neural network (TFNN) with one hidden layer and $m$ outputs is any function
$H:X\to\mathbb{R}^{m}$
of the form
$H(x)=A\,\sigma\!\bigl(T(x)\bigr)-b,$
where:
  • $T:X\to\mathbb{R}^{r}$ is a feature map of the form
    $T(x)=\bigl(w_{1}f_{1}(x)-\theta_{1},\dots,w_{r}f_{r}(x)-\theta_{r}\bigr),$
    with $r\in\mathbb{N}$, $f_{i}\in\mathcal{A}(X)$, and $w_{i},\theta_{i}\in\mathbb{R}$;
  • $A\in\mathbb{R}^{m\times r}$ and $b\in\mathbb{R}^{m}$ are the output-layer weights and bias;
  • the activation function $\sigma$ acts componentwise on vector-valued arguments.
When $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, the map $T:\mathbb{R}^{d}\to\mathbb{R}^{r}$ is an affine map. In this case, Definition 1.2 coincides exactly with the classical shallow feedforward network representation
$\Phi(x)=T_{1}\circ\sigma\circ T_{0}(x),$
where both $T_{0}$ and $T_{1}$ are affine maps.
Deep networks on topological spaces. Deep feedforward neural networks are obtained by iterating the affine–activation construction layer by layer. We adopt the same principle in the topological setting, preserving the standard compositional structure used in the classical theory.
Definition 1.3 (Deep TFNN with vector output).
Fix a topological space $X$, a basic family $\mathcal{A}(X)\subset C(X)$, an activation function $\sigma:\mathbb{R}\to\mathbb{R}$, and integers $l\geq 1$, $m\geq 1$.
A deep topological feedforward neural network (TFNN) of depth $l$ with $m$ outputs is any function
$H:X\to\mathbb{R}^{m}$
of the form
$H=T_{l+1}\circ\sigma\circ T_{l}\circ\sigma\circ\cdots\circ\sigma\circ T_{1}\circ\sigma\circ T_{0},$
where:
  • $T_{0}:X\to\mathbb{R}^{k_{0}}$ is a feature map of the form
    $T_{0}(x)=\bigl(w_{1}f_{1}(x)-\theta_{1},\dots,w_{k_{0}}f_{k_{0}}(x)-\theta_{k_{0}}\bigr),$
    with $f_{i}\in\mathcal{A}(X)$ and $w_{i},\theta_{i}\in\mathbb{R}$;
  • for $j=1,\dots,l$, $T_{j}:\mathbb{R}^{k_{j-1}}\to\mathbb{R}^{k_{j}}$ is an affine map
    $T_{j}(y)=A_{j}y-b_{j},\qquad A_{j}\in\mathbb{R}^{k_{j}\times k_{j-1}},\;b_{j}\in\mathbb{R}^{k_{j}};$
  • $T_{l+1}:\mathbb{R}^{k_{l}}\to\mathbb{R}^{m}$ is an affine output map
    $T_{l+1}(y)=A_{l+1}y-b_{l+1},\qquad A_{l+1}\in\mathbb{R}^{m\times k_{l}},\;b_{l+1}\in\mathbb{R}^{m};$
  • the activation function $\sigma$ acts componentwise on vector-valued arguments.
This definition mirrors exactly the standard representation of deep feedforward neural networks as compositions of affine maps and componentwise nonlinearities. In particular, when $X=\mathbb{R}^{d}$ and $\mathcal{A}(X)=\mathcal{L}(\mathbb{R}^{d})$, deep TFNNs reduce precisely to classical deep neural networks of the form
$\Phi(x)=T_{l+1}\circ\sigma\circ T_{l}\circ\cdots\circ\sigma\circ T_{0}(x),$
with each $T_{j}$ affine.
In this paper we are especially interested in deep narrow (or deep skinny) TFNNs, meaning deep TFNNs for which there exists a fixed integer $k\in\mathbb{N}$ such that the width of every hidden layer is at most $k$, while the depth is allowed to grow arbitrarily. In other words, in the deep narrow setting we assume that
$\sup_{j=0,\dots,l}k_{j}\leq k,$
where the bound $k$ is independent of the depth $l$.
Aim of the paper. The aim of this paper is to develop a universality theory for TFNNs with vector-valued outputs on general topological input spaces. More precisely, we seek conditions on the feature family $\mathcal{A}(X)\subset C(X)$ and on the activation function $\sigma$ under which, for every compact set $K\subset X$, deep TFNNs are dense in the space $C(K;\mathbb{R}^{m})$ of continuous $\mathbb{R}^{m}$-valued functions on $K$.
A central focus of the paper is the deep narrow setting, in which the width of each hidden layer is uniformly bounded while the depth is allowed to be arbitrary. Our goal is to identify structural assumptions ensuring that such width-constrained deep TFNNs retain universal approximation power. In this way, we extend the deep narrow universality property known for Euclidean input spaces to a broad class of non-Euclidean spaces.
Related literature. The approximation theory of neural networks is centered around the universal approximation property (UAP), also referred to as density. In the classical Euclidean setting, this property asserts that, for suitable activation functions $\sigma$, neural networks with a single hidden layer are dense in $C(K)$ for every compact set $K\subset\mathbb{R}^{d}$. The most general characterization of this phenomenon was obtained by Leshno, Lin, Pinkus, and Schocken [23], who proved that, under the assumption of continuity, an activation function $\sigma$ has the UAP if and only if it is not a polynomial. This result demonstrates the expressive power of shallow feedforward networks for a broad class of activation functions. It is worth noting that the theorem in [23] applies to a wider class of activation functions, including some that are discontinuous on sets of Lebesgue measure zero. Detailed proofs and further discussion can be found in [27, 28].
Early work on universal approximation emphasized that achieving density typically requires networks with an unbounded number of hidden neurons (see, e.g., [7, Chapter 6.4.1]). In these classical results, width was not constrained. Some studies [9, 10, 11] have shown that, for certain non-explicit but computable activation functions, universal approximation is possible even with a fixed and minimal number of hidden neurons. These results highlight that width is not always the primary source of expressive power.
Universality has also been investigated for deep architectures under explicit assumptions on the activation function. An early contribution in this direction is due to Gripenberg [8], who studied deep feedforward networks with a uniformly bounded number of neurons per hidden layer. He established universal approximation under the assumption that the activation function is continuous, nonaffine, and twice continuously differentiable in a neighborhood of some point $t$ with $\sigma^{\prime\prime}(t)\neq 0$. Subsequently, Kidger and Lyons [18] relaxed the required smoothness. They showed that universality for deep and narrow networks still holds if the activation function is continuous, nonaffine, and differentiable in a neighborhood of at least one point, with the derivative being continuous and nonzero at that point. In both works, the universality property is derived from local differentiability assumptions, without imposing any global smoothness conditions on the activation function.
In contrast, Johnson [16] obtained a negative result for a broad class of activation functions. He proved that if the activation function is uniformly continuous and can be approximated by a sequence of one-to-one functions, then deep feedforward networks with width at most $d$ in each hidden layer fail to be universal on $\mathbb{R}^{d}$, regardless of depth.
For a comprehensive and modern treatment of the mathematical analysis of deep neural networks, we refer the reader to the recent monograph by Petersen and Zech [27]. See also [12] for a focused discussion of how UAP statements are sometimes misinterpreted in the literature, and for recent results on universality under fixed architectural size, in particular fixed width and fixed depth.
The UAP for neural networks acting between infinite-dimensional spaces has also been extensively studied. In [33], the fundamentality of ridge functions in Banach spaces was established and applied to shallow networks with sigmoidal activation functions (see also [24]). Here, a subset $S$ of a topological vector space $E$ is called fundamental if the linear span of $S$ is dense in $E$. In [2], it was shown that any continuous nonlinear operator $G:V\to C(K_{2})$, defined on a compact set $V\subset C(K_{1})$, can be uniformly approximated on $V\times K_{2}$ by shallow feedforward neural networks in the following sense: for each $y\in K_{2}$, the scalar-valued map $u\mapsto G(u)(y)$ admits uniform approximation by expressions depending on finitely many point evaluations of $u$. In [25], this approach was extended to deep neural networks and led to the introduction of Deep Operator Networks (DeepONets), which are designed to approximate nonlinear continuous operators acting between spaces of continuous functions. In this setting, both the inputs and outputs of the operator are functions, and neural networks are used to learn the mapping $u\mapsto G(u)$. In [22], DeepONets were further analyzed within an encoder–decoder framework, and their approximation properties were studied for operators whose inputs lie in Hilbert spaces.
Neural networks with inputs from special topological vector spaces have also been considered. The universal approximation property for networks with inputs from Fréchet spaces and outputs in Banach spaces was established in [1]. This framework was further extended in [6], where universal approximation theorems were proved for quasi-Polish input and output spaces. Universal approximation results for functional input neural networks defined on weighted spaces, with outputs in Banach spaces, were obtained in [3].
A different perspective on neural network approximation over non-Euclidean spaces relies on abstract transfer principles, through which approximation properties established on Euclidean spaces extend to more general domains. For example, [20] investigated which modifications of the input and output layers preserve the universal approximation property. The authors showed that, under suitable conditions on topological spaces $X$ and $Y$ and continuous maps $\phi:X\to\mathbb{R}^{n}$ and $\rho:\mathbb{R}^{m}\to Y$, density of a class $\mathcal{F}\subset C(\mathbb{R}^{n},\mathbb{R}^{m})$ implies density of the composed class $\{\rho\circ f\circ\phi:f\in\mathcal{F}\}$ in $C(X,Y)$ with respect to uniform convergence on compact sets.
In [21], universal approximation results are established for neural operators (NOs) and mixtures of neural operators (MoNOs) acting between Sobolev spaces. Specifically, any nonlinear continuous operator $G^{+}:H^{s_{1}}\to H^{s_{2}}$ can be approximated with arbitrary accuracy in the $L^{2}$-norm, uniformly on compact subsets $K\subset H^{s_{1}}$, by NOs and MoNOs. In addition, [21] provides quantitative bounds on the depth, width, and rank of the approximating operators in terms of the compact set $K$ and the target accuracy.
Recent work has also established universal approximation theorems for hypercomplex-valued neural networks, including complex-, quaternion-, and Clifford-valued architectures, as well as for more general vector-valued networks (V-nets) defined over finite-dimensional algebras (see [34] and references therein).
It is worth noting that, in the shallow setting, universality for scalar-valued networks on Banach spaces of continuous functions [2] and on general topological spaces [14], as well as for networks with values in finite-dimensional algebras [34], has been established for activation functions belonging to the Tauber–Wiener class. Moreover, for topological spaces with the Hahn–Banach extension property, scalar-valued universality can be characterized by non-polynomiality of the activation function on a sufficiently small interval [13].
The present work differs from much of the above literature in two essential aspects. First, in addition to treating vector-valued shallow (and deep) networks without width constraints, we investigate deep networks with a uniform width bound and study whether universality can be retained in this setting. Second, we consider approximation of vector-valued functions defined on arbitrary topological input spaces. The only restriction of our main result is that the output space must be a finite-dimensional Euclidean space. The techniques developed here do not yet extend to neural networks with outputs in infinite-dimensional spaces. However, we hope that the results of this paper will stimulate further exploration of deep neural networks with outputs taking values in abstract infinite-dimensional spaces.
2 Universal approximation on topological spaces
In this section we formulate and prove universality results for topological feedforward neural networks whose inputs range over a general topological space $X$.
To treat approximation in the vector-valued setting, we work in the space $C(X;\mathbb{R}^{m})$ of continuous functions from $X$ into $\mathbb{R}^{m}$. In the sequel, $C(X;\mathbb{R}^{m})$ is equipped with the topology of uniform convergence on compact sets. This topology is induced by the family of seminorms
$\|g\|_{K}=\sup_{x\in K}\|g(x)\|_{\mathbb{R}^{m}},$
where $K$ ranges over all compact subsets of $X$ and $\|\cdot\|_{\mathbb{R}^{m}}$ denotes the Euclidean norm on $\mathbb{R}^{m}$ (or, equivalently, any fixed norm on $\mathbb{R}^{m}$).
A subbasis at the origin for this topology is given by the sets
$U(K,r)=\left\{g\in C(X;\mathbb{R}^{m}):\|g\|_{K}<r\right\},$
where $K\subset X$ is compact and $r>0$.
Thus, in what follows, when we say that a family of functions $\mathcal{F}$ acting from $X$ into $\mathbb{R}^{m}$ is dense in $C(X;\mathbb{R}^{m})$, we always mean density with respect to the above topology of uniform convergence on compact sets. That is, $\mathcal{F}$ is dense in $C(X;\mathbb{R}^{m})$ if, for every compact set $K\subset X$, every $g\in C(K;\mathbb{R}^{m})$, and every $\varepsilon>0$, there exists $f\in\mathcal{F}$ such that
$\|g-f\|_{K}<\varepsilon.$
Equivalently, for every compact set $K\subset X$, the collection of restrictions
$\{\,f|_{K}:f\in\mathcal{F}\,\}$
is dense in $C(K;\mathbb{R}^{m})$ with respect to the uniform seminorm $\|\cdot\|_{K}$.
Our objective is to relate approximation on the abstract input space $X$ to approximation on Euclidean spaces. This requires, on the one hand, suitable richness assumptions on the feature family $\mathcal{A}(X)\subset C(X)$, and, on the other hand, approximation properties of the activation function $\sigma$. In this context we distinguish between two cases. The first case, which covers shallow and deep TFNNs without width constraints, can be treated using a structural density condition on the feature family, called the $D$-property. The second case, corresponding to deep but narrow networks, requires additional ideas that reduce the approximation problem to a finite-dimensional setting and exploit depth to compensate for width constraints.
2.1 The $D$-property and universality without width constraints
We begin by defining a structural density property of feature families that allows approximation on $X$ to be reduced to approximation of univariate functions composed with suitable feature maps.
Definition 2.1 ($D$-property).
A family $\mathcal{A}(X)\subset C(X)$ is said to have the $D$-property if the linear span
$\mathcal{S}=\operatorname{span}\{\,u\circ f:u\in C(\mathbb{R}),\ f\in\mathcal{A}(X)\,\}$ (2.1)
is dense in $C(X)$ with respect to the topology of uniform convergence on compact sets.
Equivalently, $\mathcal{A}(X)$ has the $D$-property if and only if for every compact set $K\subset X$, every function $g\in C(K)$, and every $\varepsilon>0$, there exist an integer $m\in\mathbb{N}$, feature maps $f_{1},\dots,f_{m}\in\mathcal{A}(X)$, and univariate continuous functions $u_{1},\dots,u_{m}\in C(\mathbb{R})$ such that
$\max_{x\in K}\left|g(x)-\sum_{i=1}^{m}u_{i}\bigl(f_{i}(x)\bigr)\right|<\varepsilon.$
Remark 2.1.
The $D$-property asserts that the feature family $\mathcal{A}(X)$ is sufficiently rich to approximate arbitrary continuous functions on $X$ through finite superpositions with continuous univariate functions. In particular, the $D$-property implies a strong separation capability: for any compact set $K\subset X$ and any two distinct points $x,y\in K$, there exists a function in the span (2.1) that takes different values at $x$ and $y$.
In the classical setting $X=\mathbb{R}^{d}$, the role of $\mathcal{A}(X)$ is played by the family of linear functionals $\mathbf{x}\mapsto\mathbf{a}\cdot\mathbf{x}$. In this case, density of finite superpositions of ridge functions $u(\mathbf{a}\cdot\mathbf{x})$ is a well-studied topic in approximation theory (see, e.g., [11, 29]).
2.2 A univariate approximation assumption
To realize compositions of the form $u\circ f$ by neural networks, we require that the activation function be able to approximate arbitrary continuous functions of a single real variable. The following abstract assumption captures this requirement.
Definition 2.2 (Univariate UAP for $\sigma$).
An activation function $\sigma:\mathbb{R}\to\mathbb{R}$ is said to satisfy the univariate universal approximation property if for every compact interval $a,b]\subset\mathbb{R}$ and every $\varepsilon>0$, each function $u\in C([a,b])$ can be approximated uniformly on $[a,b]$ by finite linear combinations of shifted and rescaled copies of $\sigma$, that is, by functions of the form
$t\mapsto\sum_{j=1}^{N}c_{j}\,\sigma(w_{j}t-\theta_{j}),\qquad c_{j},w_{j},\theta_{j}\in\mathbb{R}.$
Remark 2.2.
Definition 2.2 concerns approximation by shallow (single-hidden-layer) univariate neural networks. This property is classical and is satisfied by many commonly used activation functions. In particular, it holds for continuous non-polynomial activations [23] and characterizes the Tauber–Wiener class considered in several papers (see, e.g., [2, 14, 34]). In the present section, only this univariate approximation property is required.
2.3 Universality of TFNNs without width constraints
We now establish a universality statement for topological feedforward neural networks with vector-valued outputs. Throughout this subsection, no restriction is imposed on the width: the number of hidden neurons depends both on the target function and on the desired accuracy.
Theorem 2.1 (Universality from the $D$-property).
Let $X$ be a topological space and let $\mathcal{A}(X)\subset C(X)$ be a feature family with the $D$-property. Assume that the activation function $\sigma:\mathbb{R}\to\mathbb{R}$ satisfies the univariate universal approximation property of Definition 2.2. Then for every compact set $K\subset X$, every function $g\in C(K;\mathbb{R}^{m})$, and every $\varepsilon>0$, there exists a TFNN $H\in\mathcal{N}_{1}^{(m)}(\sigma)$ such that
$\|g-H\|_{K}<\varepsilon.$
Consequently, for every $l\geq 1$, the deep TFNN class $\mathcal{N}_{l}^{(m)}(\sigma)$ is dense in $C(X;\mathbb{R}^{m})$.
Proof.
Fix a compact set $K\subset X$, a vector-valued function
$g=(g_{1},\dots,g_{m})\in C(K;\mathbb{R}^{m}),$
and $\varepsilon>0$. We begin by constructing, for $l=1$, a TFNN $H=(H_{1},\dots,H_{m})$ satisfying
$\|g-H\|_{K}=\sup_{x\in K}\|g(x)-H(x)\|_{\mathbb{R}^{m}}<\varepsilon.$
With the Euclidean norm on $\mathbb{R}^{m}$,
$\|g-H\|_{K}=\sup_{x\in K}\left(\sum_{k=1}^{m}|g_{k}(x)-H_{k}(x)|^{2}\right)^{1/2}\leq\left(\sum_{k=1}^{m}\|g_{k}-H_{k}\|_{K}^{2}\right)^{1/2},$
where for scalar-valued functions we use the seminorm $\|h\|_{K}=\sup_{x\in K}|h(x)|$.
Accordingly, we proceed by constructing, for each component $k=1,\dots,m$, a scalar TFNN $H_{k}$ such that
$\|g_{k}-H_{k}\|_{K}<\frac{\varepsilon}{\sqrt{m}}.$
Fix $k\in\{1,\dots,m\}$. The function $g_{k}$ is continuous on the compact set $K$. By the $D$-property of $\mathcal{A}(X)$, the linear span of compositions
$\operatorname{span}\{u\circ f:\ u\in C(\mathbb{R}),\ f\in\mathcal{A}(X)\}$
is dense in $C(X)$ in the topology of uniform convergence on compact sets. Applying this density to the restriction on $K$, we can approximate $g_{k}$ uniformly on $K$ by a finite sum of such compositions. Concretely, there exist an integer $M_{k}\in\mathbb{N}$, feature maps $f_{k,1},\dots,f_{k,M_{k}}\in\mathcal{A}(X)$, and continuous functions $u_{k,1},\dots,u_{k,M_{k}}\in C(\mathbb{R})$ such that
$\sup_{x\in K}\left|g_{k}(x)-\sum_{i=1}^{M_{k}}u_{k,i}\bigl(f_{k,i}(x)\bigr)\right|<\frac{\varepsilon}{2\sqrt{m}}.$ (2.2)
Fix $i\in\{1,\dots,M_{k}\}$. Since $f_{k,i}$ is continuous and $K$ is compact, the image
$f_{k,i}(K)\subset\mathbb{R}$
is compact. Hence there exists a compact interval $a_{k,i},b_{k,i}]$ with
$f_{k,i}(K)\subset[a_{k,i},b_{k,i}].$
(For example, take $a_{k,i}=\min_{x\in K}f_{k,i}(x)$ and $b_{k,i}=\max_{x\in K}f_{k,i}(x)$.)
Thus, to approximate $u_{k,i}(f_{k,i}(x))$ uniformly on $x\in K$, it is enough to approximate the scalar function $u_{k,i}$ uniformly on the interval $a_{k,i},b_{k,i}]$.
By Definition 2.2, for each $i$ there exist an integer $N_{k,i}\in\mathbb{N}$ and real parameters $c_{k,i,j},w_{k,i,j},\theta_{k,i,j}\in\mathbb{R}$ such that
$\sup_{t\in[a_{k,i},b_{k,i}]}\left|u_{k,i}(t)-\sum_{j=1}^{N_{k,i}}c_{k,i,j}\,\sigma(w_{k,i,j}t-\theta_{k,i,j})\right|<\frac{\varepsilon}{2\sqrt{m}\,M_{k}}.$ (2.3)
Because $f_{k,i}(K)\subset[a_{k,i},b_{k,i}]$, the same estimate holds when we substitute $t=f_{k,i}(x)$ with $x\in K$. Substituting $t=f_{k,i}(x)$ in (2.3) gives
$\sup_{x\in K}\left|u_{k,i}\bigl(f_{k,i}(x)\bigr)-\sum_{j=1}^{N_{k,i}}c_{k,i,j}\,\sigma\bigl(w_{k,i,j}f_{k,i}(x)-\theta_{k,i,j}\bigr)\right|<\frac{\varepsilon}{2\sqrt{m}\,M_{k}}.$
Now sum over $i=1,\dots,M_{k}$ and use the triangle inequality:
$\sup_{x\in K}\left|\sum_{i=1}^{M_{k}}u_{k,i}\bigl(f_{k,i}(x)\bigr)-\sum_{i=1}^{M_{k}}\sum_{j=1}^{N_{k,i}}c_{k,i,j}\,\sigma\bigl(w_{k,i,j}f_{k,i}(x)-\theta_{k,i,j}\bigr)\right|$
$\qquad\leq\sum_{i=1}^{M_{k}}\sup_{x\in K}\left|u_{k,i}\bigl(f_{k,i}(x)\bigr)-\sum_{j=1}^{N_{k,i}}c_{k,i,j}\,\sigma\bigl(w_{k,i,j}f_{k,i}(x)-\theta_{k,i,j}\bigr)\right|<\frac{\varepsilon}{2\sqrt{m}}.$ (2.4)
Thus, the finite sum of compositions in (2.2) is uniformly approximated on $K$ by a finite linear combination of terms of the form $\sigma(wf(x)-\theta)$.
Define the scalar function
$H_{k}(x)=\sum_{i=1}^{M_{k}}\sum_{j=1}^{N_{k,i}}c_{k,i,j}\,\sigma\bigl(w_{k,i,j}f_{k,i}(x)-\theta_{k,i,j}\bigr),\qquad x\in X.$
Then $H_{k}\in\mathcal{N}_{1}^{(1)}(\sigma)$, which is built from feature maps in $\mathcal{A}(X)$ and activation $\sigma$.
For each $x\in K$, we insert and subtract the intermediate approximation from (2.2) and use the triangle inequality:
$|g_{k}(x)-H_{k}(x)|$ $\leq\left|g_{k}(x)-\sum_{i=1}^{M_{k}}u_{k,i}\bigl(f_{k,i}(x)\bigr)\right|+\left|\sum_{i=1}^{M_{k}}u_{k,i}\bigl(f_{k,i}(x)\bigr)-H_{k}(x)\right|$
$<\frac{\varepsilon}{2\sqrt{m}}+\frac{\varepsilon}{2\sqrt{m}}=\frac{\varepsilon}{\sqrt{m}},$
where we used (2.2) for the first term and (2.4) for the second term. Taking the supremum over $x\in K$ in all three absolute values yields
$\|g_{k}-H_{k}\|_{K}<\frac{\varepsilon}{\sqrt{m}}.$
We repeat the above for each component $k=1,\dots,m$, obtaining scalar approximants $H_{1},\dots,H_{m}$. Define
$H(x)=(H_{1}(x),\dots,H_{m}(x)),\qquad x\in X.$
By construction, $H$ is a single hidden layer TFNN with $m$ outputs, i.e. $H\in\mathcal{N}_{1}^{(m)}(\sigma)$.
Finally, since $\|g_{k}-H_{k}\|_{K}<\varepsilon/\sqrt{m}$ for each $k=1,\dots,m$, it follows that
$\|g-H\|_{K}=\sup_{x\in K}\left(\sum_{k=1}^{m}|g_{k}(x)-H_{k}(x)|^{2}\right)^{1/2}\leq\left(\sum_{k=1}^{m}\|g_{k}-H_{k}\|_{K}^{2}\right)^{1/2}<\left(\sum_{k=1}^{m}\frac{\varepsilon^{2}}{m}\right)^{1/2}=\varepsilon.$
For the deep classes, note that every shallow network may be viewed as a deep network by inserting identity (or otherwise trivial) affine layers. Formally, by Definition 1.3, we have
$\mathcal{N}_{1}^{(m)}(\sigma)\subset\mathcal{N}_{l}^{(m)}(\sigma),\qquad l\geq 1.$
If a set $B$ of a topological vector space is dense, then any superset of $B$ is also dense. Hence each $\mathcal{N}_{l}^{(m)}(\sigma)$ is dense as well. This completes the proof. ∎
2.4 Locally convex spaces
We now analyze an important and widely applicable special case: locally convex topological vector spaces. In this setting there is a natural choice of feature family, namely the continuous dual space.
Throughout this subsection, all locally convex topological vector spaces are assumed to be Hausdorff.
Let $X$ be a locally convex topological vector space over $\mathbb{R}$, and denote by $X^{*}$ its continuous dual. Elements of $X^{*}$ play the role of linear measurements of the input, generalizing the inner products $\mathbf{w}\cdot\mathbf{x}$ that appear in classical neural networks on $\mathbb{R}^{d}$.
Theorem 2.2 (Universality on locally convex spaces).
Let $X$ be a locally convex topological vector space and let $\sigma:\mathbb{R}\to\mathbb{R}$ satisfy the univariate universal approximation property of Definition 2.2. Then, for every integer $m\geq 1$ and every compact set $K\subset X$, the TFNN class $\mathcal{N}_{1}^{(m)}(\sigma)$ associated with the feature family $\mathcal{A}(X)=X^{*}$ is dense in $C(K;\mathbb{R}^{m})$ with respect to the uniform norm on $K$. Consequently, for every depth $l\geq 1$, the deep TFNN class $\mathcal{N}_{l}^{(m)}(\sigma)$ is dense in $C(K;\mathbb{R}^{m})$.
Proof.
We first verify that the feature family $\mathcal{A}(X)=X^{*}$ satisfies the $D$-property. Define
$\mathcal{S}=\operatorname{span}\{\,u\circ f:u\in C(\mathbb{R}),\ f\in X^{*}\,\}\subset C(X).$
We show that $\mathcal{S}$ is dense in $C(X)$ with respect to uniform convergence on compact sets.
Let $K\subset X$ be an arbitrary compact set. Consider the set
$\mathcal{E}=\operatorname{span}\{\,e^{f(x)}:f\in X^{*}\,\}\subset C(X).$
This set is an algebra, since
$e^{f_{1}(x)}e^{f_{2}(x)}=e^{(f_{1}+f_{2})(x)}\quad\text{and}\quad f_{1}+f_{2}\in X^{*},$
and it contains the constant functions because $e^{0}=1$.
Moreover, $\mathcal{E}$ separates points of $K$. Indeed, since $X$ is locally convex, the Hahn–Banach extension theorem holds. It is a consequence of this theorem that for any $x,y\in K$ with $x\neq y$ there exists $f\in X^{*}$ such that $f(x)\neq f(y)$ (see, e.g., [31, Theorem 3.6]). Consequently,
$e^{f(x)}\neq e^{f(y)}.$
By the Stone–Weierstrass theorem [32], the restriction $\mathcal{E}|_{K}$ is dense in $C(K)$ with respect to the uniform norm. On the other hand, each generator $e^{f(x)}$ belongs to $\mathcal{S}$ (take $u(t)=e^{t}$), hence $\mathcal{E}\subset\mathcal{S}$ and therefore $\mathcal{E}|_{K}\subset\mathcal{S}|_{K}$. It follows that $\mathcal{S}|_{K}$ is dense in $C(K)$. As $K\subset X$ was arbitrary, this proves that $\mathcal{A}(X)=X^{*}$ has the $D$-property.
We now apply Theorem 2.1. Let $m\geq 1$, let $K\subset X$ be compact, and let $g\in C(K;\mathbb{R}^{m})$ be arbitrary. Since $\mathcal{A}(X)=X^{*}$ has the $D$-property and $\sigma$ satisfies the univariate universal approximation property, Theorem 2.1 implies that for every $\varepsilon>0$ there exists a TFNN
$H\in\mathcal{N}_{1}^{(m)}(\sigma)$
such that
$\|g-H\|_{K}<\varepsilon.$
Finally, since every shallow TFNN can be regarded as a deep TFNN by inserting additional affine and activation layers, we have
$\mathcal{N}_{1}^{(m)}(\sigma)\subset\mathcal{N}_{l}^{(m)}(\sigma)\quad\text{for all }l\geq 1.$
Hence density holds for all depths. ∎
A classical and influential result in this direction is due to Chen and Chen [2, Theorem 4], who studied approximation of continuous functionals by shallow neural networks on a compact subset $V\subset C(Y)$. Their theorem shows that, if the activation function belongs to the Tauber–Wiener class, then every continuous functional on $V$ can be uniformly approximated by shallow neural network functionals whose inputs depend only on finitely many point evaluations $u(x_{1}),\dots,u(x_{k})$, where $u\in V$.
In the present framework, a corresponding approximation result can be obtained as a consequence of the universality theorem for locally convex spaces (Theorem 2.2), together with the fact that, on compact subsets of $C(Y)$, continuous linear functionals can be uniformly approximated by finite linear combinations of point-evaluation functionals. The following theorem makes this connection explicit for continuous activation functions.
Theorem 2.3 (Chen and Chen [2]).
Let $Y$ be a compact metric space and let $V\subset C(Y)$ be compact. Assume that the activation $\sigma\in C(\mathbb{R})$ satisfies the univariate universal approximation property of Definition 2.2. Then for every continuous functional $f\in C(V)$ and every $\varepsilon>0$ there exist integers $N,k\geq 1$, points $x_{1},\dots,x_{k}\in K$, and real parameters $c_{i},\theta_{i},\xi_{ij}$ ($i=1,\dots,N$, $j=1,\dots,k$) such that
$\left|f(u)-\sum_{i=1}^{N}c_{i}\,\sigma\!\left(\sum_{j=1}^{k}\xi_{ij}\,u(x_{j})-\theta_{i}\right)\right|<\varepsilon$
holds for all $u\in V$.
Proof.
We view $C(Y)$ as a Banach space and hence as a Hausdorff locally convex space. Apply Theorem 2.2 with $X=C(Y)$, compact set $V\subset X$, and $m=1$. Then there exist $N\geq 1$, functionals $\ell_{1},\dots,\ell_{N}\in C(Y)^{*}$, and parameters $c_{i},\theta_{i}\in\mathbb{R}$ such that
$\sup_{u\in V}\left|f(u)-\sum_{i=1}^{N}c_{i}\,\sigma(\ell_{i}(u)-\theta_{i})\right|<\frac{\varepsilon}{2}.$ (2.5)
Next we approximate each $\ell_{i}$ uniformly on $V$ by finitely many point evaluations. Since $V\subset C(Y)$ is compact, it is equicontinuous by the Arzelà–Ascoli theorem. Let $\ell\in C(Y)^{*}$ be arbitrary. By the Riesz representation theorem, $\ell(u)=\int_{Y}u\,d\mu$ for a finite signed Borel measure $\mu$ on $Y$. Equicontinuity of $V$ implies that for every $\delta>0$ there exist points $x_{1},\dots,x_{k}\in Y$ and coefficients $\alpha_{1},\dots,\alpha_{k}\in\mathbb{R}$ such that
$\sup_{u\in V}\left|\ell(u)-\sum_{j=1}^{k}\alpha_{j}\,u(x_{j})\right|<\delta.$ (2.6)
Now apply (2.6) to each $\ell_{i}$, $i=1,\dots,N$. For each $i$ this yields a finite set of sample points $\{x_{i,1},\dots,x_{i,k_{i}}\}\subset Y$ and coefficients $\alpha_{i,j}$ such that
$\sup_{u\in V}\left|\ell_{i}(u)-\sum_{j=1}^{k_{i}}\alpha_{i,j}\,u(x_{i,j})\right|<\delta.$
Let $\{x_{1},\dots,x_{k}\}$ be the union of all these sampling points over $i=1,\dots,N$. By setting the coefficients to zero if necessary, each sum may be rewritten over this common set, yielding coefficients $\xi_{ij}$ such that
$\sup_{u\in V}\left|\ell_{i}(u)-\sum_{j=1}^{k}\xi_{ij}\,u(x_{j})\right|<\delta,\qquad i=1,\dots,N.$
Since $V$ is compact and each $\ell_{i}$ is continuous, the set $\ell_{i}(V)\subset\mathbb{R}$ is compact. Therefore $\sigma$ is uniformly continuous on a compact interval containing both $\ell_{i}(V)$ and the approximants $\sum_{j}\xi_{ij}u(x_{j})$ for $u\in V$ (for $\delta$ small enough). Choose $\delta>0$ so that
$|s-t|<\delta\quad\Longrightarrow\quad|\sigma(s-\theta_{i})-\sigma(t-\theta_{i})|<\frac{\varepsilon}{2\,\sum_{i=1}^{N}|c_{i}|+1}\qquad(i=1,\dots,N).$
Then for every $u\in V$,
$\left|c_{i}\sigma(\ell_{i}(u)-\theta_{i})-c_{i}\sigma\!\left(\sum_{j=1}^{k}\xi_{ij}u(x_{j})-\theta_{i}\right)\right|\leq|c_{i}|\cdot\frac{\varepsilon}{2\,\sum_{i=1}^{N}|c_{i}|+1}.$
Summing over $i=1,\dots,N$ gives
$\sup_{u\in V}\left|\sum_{i=1}^{N}c_{i}\,\sigma(\ell_{i}(u)-\theta_{i})-\sum_{i=1}^{N}c_{i}\,\sigma\!\left(\sum_{j=1}^{k}\xi_{ij}u(x_{j})-\theta_{i}\right)\right|<\frac{\varepsilon}{2}.$ (2.7)
Finally, combine (2.5) and (2.7) by the triangle inequality. This yields the desired representation with error $\varepsilon$. ∎
Remark 2.3.
Theorem 2.2 shows that locally convex spaces (including Banach spaces and Fréchet spaces) form a broad and natural class of input domains for which universality of topological feedforward neural networks holds. This universality result is driven by the abstract $D$-property and relies on the separating power of the continuous dual, rather than on any finite-dimensional structure of the input space. The subsequent Theorem 2.3 of Chen and Chen may be viewed as a special case of this general principle applied to the Banach space $C(Y)$. In the next section, we turn to the substantially more delicate question of whether universality can be retained under uniform width constraints, that is, in the deep narrow framework.
3 Universality of deep narrow TFNNs
In this section we address the question announced in the previous remark: whether universal approximation on a general topological space $X$ can be achieved by deep networks of uniformly bounded width. As discussed earlier, this problem is more delicate than universality without width constraints. The $D$-property yields approximation by sums of many univariate compositions, but it does not by itself control how many neurons must appear in each hidden layer.
Our strategy is to impose an additional finite-dimensional composition assumption on the feature family $\mathcal{A}(X)$. Under this assumption, approximation on $X$ reduces to approximation on a compact subset of $\mathbb{R}^{n}$. We can then invoke the deep narrow universality theorem of Kidger and Lyons [18] for Euclidean inputs and vector-valued outputs.
3.1 Finite-dimensional composition via feature maps
We begin by formalizing the composition property. The key point, compared with the $D$-property, is that for each compact set $K\subset X$ and each output dimension $m$ one fixes a finite collection of feature maps, and approximation of different target functions is achieved solely by varying the outer function.
Definition 3.1 (Finite-dimensional composition).
A feature family $\mathcal{A}(X)\subset C(X)$ is said to have the finite-dimensional composition property of order $n$ if for every compact set $K\subset X$ and every integer $m\geq 1$ there exist functions $f_{1},\dots,f_{n}\in\mathcal{A}(X)$ such that, with $F=(f_{1},\dots,f_{n}):X\to\mathbb{R}^{n}$, the set
$\{\,u\circ F|_{K}:u\in C(\mathbb{R}^{n};\mathbb{R}^{m})\,\}$
is dense in $C(K;\mathbb{R}^{m})$.
Remark 3.1.
Definition 3.1 means that, on each compact subset $K$ of $X$, every continuous $\mathbb{R}^{m}$-valued function can be approximated arbitrarily well by first mapping $x\in X$ to the Euclidean feature vector $F(x)\in\mathbb{R}^{n}$ and then applying a continuous map $u:\mathbb{R}^{n}\to\mathbb{R}^{m}$. In contrast to the $D$-property, the same finite set of features $f_{1},\dots,f_{n}$ is used for all target functions and all accuracies; only the mapping $u$ is allowed to change.
The finite-dimensional composition property of Definition 3.1 is formulated as a density condition. In many situations, however, this property holds in a stronger form, namely with exact representations rather than mere approximation. The following proposition identifies situations in which such a strong finite-dimensional composition property holds on compact sets.
Proposition 3.1 (Strong composition property).
Let $X$ be a topological space and let $K\subset X$ be compact.
(i) Suppose that there exists a feature map
$F=(f_{1},\dots,f_{n}):X\to\mathbb{R}^{n},\qquad f_{i}\in\mathcal{A}(X),$
such that the restriction $F|_{K}:K\to F(K)$ is a topological embedding. Then the finite-dimensional composition property on $K$ holds in a strong form: for every integer $m\geq 1$ and every function $g\in C(K;\mathbb{R}^{m})$ there exists a function $u\in C(\mathbb{R}^{n};\mathbb{R}^{m})$ such that
$g=u\circ F|_{K}.$
(ii) Suppose that there exists a topological embedding
$E=(E_{1},\dots,E_{n}):K\hookrightarrow\mathbb{R}^{n}$
and functions $f_{1},\dots,f_{n}\in\mathcal{A}(X)$ such that
$f_{i}|_{K}=E_{i},\qquad i=1,\dots,n.$
Let $F=(f_{1},\dots,f_{n}):X\to\mathbb{R}^{n}$. Then the same strong finite-dimensional composition property on $K$ holds: for every $m\geq 1$ and every $g\in C(K;\mathbb{R}^{m})$ there exists $u\in C(\mathbb{R}^{n};\mathbb{R}^{m})$ such that
$g=u\circ F|_{K}.$
Proof.
(i) Since $K$ is compact and $F$ is continuous, the set $F(K)\subset\mathbb{R}^{n}$ is compact and hence closed. As $F|_{K}:K\to F(K)$ is a topological embedding, it is a homeomorphism between $K$ and $F(K)$. Given any $g\in C(K;\mathbb{R}^{m})$, define a function $v:F(K)\to\mathbb{R}^{m}$ by
$v(y)=g\bigl((F|_{K})^{-1}(y)\bigr),\qquad y\in F(K).$
Then $v$ is continuous on the compact set $F(K)$. By the Tietze extension theorem [35, Theorem 15.8] (applied componentwise), there exists a function $u\in C(\mathbb{R}^{n};\mathbb{R}^{m})$ such that $u|_{F(K)}=v$. Consequently, for all $x\in K$,
$(u\circ F)(x)=u(F(x))=v(F(x))=g(x),$
and hence $g=u\circ F|_{K}$.
(ii) By assumption, $E:K\to\mathbb{R}^{n}$ is a topological embedding and
$F|_{K}=(f_{1}|_{K},\dots,f_{n}|_{K})=(E_{1},\dots,E_{n})=E.$
Hence $F|_{K}$ is a topological embedding as well, and the conclusion follows directly from part (i). ∎
Remark 3.2.
Proposition 3.1 separates two distinct issues. Part (i) shows that, once a finite-dimensional feature map $F=(f_{1},\dots,f_{n})$ with $f_{i}\in\mathcal{A}(X)$ is available whose restriction to a compact set $K$ is a topological embedding, the finite-dimensional composition property on $K$ holds in a strong form, in fact yielding exact representations $g=u\circ F|_{K}$ for all $g\in C(K;\mathbb{R}^{m})$.
Part (ii) clarifies how such feature maps may arise in practice. In many settings, the existence of an embedding $E:K\hookrightarrow\mathbb{R}^{n}$ is a purely topological fact, independent of the choice of feature family $\mathcal{A}(X)$ (for example, by classical embedding theorems for compact metric spaces). The additional requirement in part (ii) is that the coordinate functions of such an embedding can be realized by elements of $\mathcal{A}(X)$. In this case, the embedding $E$ can be extended to a feature map $F$ on $X$, and the strong finite-dimensional composition property on $K$ follows.
3.2 Deep narrow TFNNs and a width constraint
We now introduce the width-constrained class that will be used in the sequel.
Fix $n\in\mathbb{N}$ and a feature map $F:X\to\mathbb{R}^{n}$ of the form $F=(f_{1},\dots,f_{n})$ with $f_{i}\in\mathcal{A}(X)$. Any (classical) deep feedforward network $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with activation $\sigma$ then induces a vector-valued network on $X$ by composition:
$H(x)=\Phi(F(x)),\qquad x\in X.$
Intuitively, $F$ provides the input coordinates, and $\Phi$ performs the standard Euclidean deep processing. We use the standard notion of width for Euclidean deep networks: the width of a network is the maximum number of neurons in any hidden layer of the Euclidean network $\Phi$.
Definition 3.2 (Deep narrow TFNNs based on $F$).
Let $k\in\mathbb{N}$ and $m\geq 1$. We denote by $\mathcal{N}^{(k,m)}(X,\sigma;F)$ the class of all functions $H:X\to\mathbb{R}^{m}$ for which there exists a (classical) deep feedforward network $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with activation function $\sigma$ and with width at most $k$ such that
$H(x)=\Phi(F(x)),\qquad x\in X.$
We refer to the elements of $\mathcal{N}^{(k,m)}(X,\sigma;F)$ as deep narrow TFNNs with $m$ outputs based on $F$.
3.3 Universality theorem under a width bound
We now state the universality theorem for deep narrow TFNNs.
Theorem 3.1 (Universality of deep narrow TFNNs).
Let $X$ be a topological space and let $\mathcal{A}(X)\subset C(X)$ satisfy the finite-dimensional composition property of order $n$ (Definition 3.1). Assume that $\sigma:\mathbb{R}\to\mathbb{R}$ is continuous and nonaffine, and that there exists a point $t_{0}\in\mathbb{R}$ and an open neighborhood $U$ of $t_{0}$ such that $\sigma$ is differentiable on $U$ and $\sigma^{\prime}$ is continuous and nonzero at $t_{0}$.
Then the following holds: for every integer $m\geq 1$ and every compact set $K\subset X$, if $f_{1},\dots,f_{n}\in\mathcal{A}(X)$ are feature maps provided by Definition 3.1 for this $K$ and $m$, and $F=(f_{1},\dots,f_{n}):X\to\mathbb{R}^{n}$, then for every function $g\in C(K;\mathbb{R}^{m})$ and every $\varepsilon>0$, there exists a function
$H\in\mathcal{N}^{(n+m+2,m)}(X,\sigma;F)$
such that
$\|g-H\|_{K}<\varepsilon.$
That is, the class $\mathcal{N}^{(n+m+2,m)}(X,\sigma;F)|_{K}$ is dense in $C(K;\mathbb{R}^{m})$.
Proof.
Fix $m\geq 1$, a compact set $K\subset X$, a target function $g\in C(K;\mathbb{R}^{m})$, and $\varepsilon>0$.
By the finite-dimensional composition property (Definition 3.1) applied to this particular compact set $K$ and output dimension $m$, there exist feature maps $f_{1},\dots,f_{n}\in\mathcal{A}(X)$ and a continuous map $u\in C(\mathbb{R}^{n};\mathbb{R}^{m})$ such that, writing
$F=(f_{1},\dots,f_{n}):X\to\mathbb{R}^{n},$
we have
$\|g-u\circ F\|_{K}<\frac{\varepsilon}{2}.$ (3.1)
Here the feature map $F$ depends on $K$ and $m$, while the map $u$ additionally depends on $g$ and $\varepsilon$.
Since $F$ is continuous and $K$ is compact, the image
$K_{F}:=F(K)\subset\mathbb{R}^{n}$
is compact. Consider the restriction $u|_{K_{F}}\in C(K_{F};\mathbb{R}^{m})$.
Under the stated regularity assumptions on $\sigma$, the deep narrow universality theorem of Kidger and Lyons [18, Theorem 3.2] yields the following conclusion: for the compact set $K_{F}\subset\mathbb{R}^{n}$ and tolerance $\varepsilon/2$, there exists a classical deep feedforward network
$\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$
with activation function $\sigma$ and with width at most $n+m+2$ such that
$\max_{y\in K_{F}}\|u(y)-\Phi(y)\|_{\mathbb{R}^{m}}<\frac{\varepsilon}{2}.$ (3.2)
Define
$H(x)=\Phi(F(x)),\qquad x\in X.$
By Definition 3.2, since $\Phi$ has width at most $n+m+2$, we have
$H\in\mathcal{N}^{(n+m+2,m)}(X,\sigma;F).$
Let $x\in K$. Then $F(x)\in K_{F}$, and therefore (3.2) implies
$\|u(F(x))-\Phi(F(x))\|_{\mathbb{R}^{m}}<\frac{\varepsilon}{2}.$
Together with (3.1), we obtain
$\|g(x)-H(x)\|_{\mathbb{R}^{m}}\leq\|g(x)-u(F(x))\|_{\mathbb{R}^{m}}+\|u(F(x))-\Phi(F(x))\|_{\mathbb{R}^{m}}<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.$
Taking the maximum over $x\in K$ yields $\|g-H\|_{K}<\varepsilon$, which completes the proof. ∎
Remark 3.3 (Scope and realizability of Theorem 3.1).
The finite-dimensional composition property of order $n$ imposes a structural restriction on the input space $X$. It requires that for every compact set $K\subset X$ and every output dimension $m$, there exist continuous feature maps
$F=(f_{1},\dots,f_{n}):X\to\mathbb{R}^{n}$
such that every continuous $\mathbb{R}^{m}$-valued function on $K$ can be uniformly approximated by functions of the form $u\circ F$ with $u\in C(\mathbb{R}^{n};\mathbb{R}^{m})$.
In particular, for each compact set $K\subset X$, the restriction $F|_{K}$ must be injective. Indeed, if $x,y\in K$ with $x\neq y$ and $F(x)=F(y)$, then
$(u\circ F)(x)=(u\circ F)(y)\quad\text{for every }u\in C(\mathbb{R}^{n};\mathbb{R}^{m}).$
Assume for the moment that $X$ is Hausdorff. Then $K$ is compact Hausdorff and hence normal. Since $\{x\}$ and $\{y\}$ are disjoint closed subsets of $K$, Urysohn’s lemma [35, Theorem 15.6] yields a function $h\in C(K)$ such that $h(x)\neq h(y)$. Consequently, $g=(h,0,\dots,0)\in C(K;\mathbb{R}^{m})$ satisfies $g(x)\neq g(y)$. Such a function cannot be approximated uniformly on $K$ by functions of the form $u\circ F$, contradicting Definition 3.1. Hence $F|_{K}$ is injective.
We see that if $X$ is Hausdorff, then $F|_{K}$ is a continuous injective map from the compact space $K$ into the Hausdorff space $\mathbb{R}^{n}$, and therefore a topological embedding of $K$ into $\mathbb{R}^{n}$ (a homeomorphism onto its image $F(K)$ [5, Theorem 3.1.13]). On the other hand, whenever $F|_{K}$ is a topological embedding, Proposition 3.1(i) implies that every function $g\in C(K;\mathbb{R}^{m})$ admits an exact representation
$g=u\circ F|_{K}\qquad\text{for some }u\in C(\mathbb{R}^{n};\mathbb{R}^{m}).$
It follows that, when $X$ is Hausdorff, the finite-dimensional composition property on a compact set $K\subset X$ automatically implies its strong form on $K$.
In the class of separable metric spaces, the above composition property is closely related to finiteness of the topological (covering) dimension. More precisely, if $X$ is a separable metric space with $\dim_{\mathrm{top}}X\leq d$, then by the Menger–Nöbeling embedding theorem [4, Theorem 1.11.4] there exists a topological embedding
$E:X\hookrightarrow\mathbb{R}^{2d+1}.$
If, in addition, the feature family $\mathcal{A}(X)$ is sufficiently rich to contain the coordinate functions of such an embedding, that is, if
$E=(E_{1},\dots,E_{2d+1})\quad\text{with}\quad E_{i}\in\mathcal{A}(X),$
then Definition 3.1 is satisfied with $n=2d+1$ by taking $f_{i}=E_{i}$ and $F=E$. In this case, the same fixed feature map $F$ embeds every compact subset of $X$ into $\mathbb{R}^{2d+1}$, and Proposition 3.1(ii) yields exact representations on each compact set.
As a special case, when $X=\mathbb{R}^{n}$ and $\mathcal{A}(X)$ consists of the coordinate projections, the feature map $F$ coincides with the identity on $\mathbb{R}^{n}$. In this setting, Theorem 3.1 reduces exactly to the deep narrow universality theorem of Kidger and Lyons for Euclidean input spaces and vector-valued outputs.
3.4 Deep narrow universality based on Ostrand inner functions
We begin by recalling the classical Kolmogorov superposition theorem (KST) and its extension due to Ostrand. For background, refinements, and generalizations of KST we refer to [17, Chapter 1] and [11, Chapter 4]. Kolmogorov’s original theorem [19] asserts that for the unit cube $\mathbb{I}^{d}$, $\mathbb{I}=[0,1]$, $d\geq 2$, there exist continuous functions $s_{q}\in C(\mathbb{I}^{d})$, $q=1,\dots,2d+1$, of the form
$s_{q}(x_{1},\dots,x_{d})=\sum_{p=1}^{d}\varphi_{pq}(x_{p}),\qquad\varphi_{pq}\in C(\mathbb{I}),$
such that every $f\in C(\mathbb{I}^{d})$ admits a representation
$f(x)=\sum_{q=1}^{2d+1}g_{q}\bigl(s_{q}(x)\bigr),\qquad g_{q}\in C(\mathbb{R}).$
This deep result, which solved Hilbert’s 13th problem in the negative, has stimulated extensive research and has also been discussed at length in connection with neural networks (see, e.g., [15] and the references therein).
Ostrand [26] extended KST to general compact metric spaces as follows.
Theorem 3.2 (Ostrand [26]).
For $p=1,\dots,n_{0}$, let $X_{p}$ be a compact metric space of finite (topological) dimension $d_{p}$, and set $M=\sum_{p=1}^{n_{0}}d_{p}$. There exist universal continuous functions $\psi_{pq}:X_{p}\to[0,1]$, $p=1,\dots,n_{0}$, $q=1,\dots,2M+1$, such that every $f\in C\!\left(\prod_{p=1}^{n_{0}}X_{p}\right)$ is representable in the form
$f(x_{1},\dots,x_{n_{0}})=\sum_{q=1}^{2M+1}h_{q}\!\left(\sum_{p=1}^{n_{0}}\psi_{pq}(x_{p})\right),\qquad h_{q}\in C(\mathbb{R}).$
Define the associated Ostrand inner functions on $X=\prod_{p=1}^{n_{0}}X_{p}$ by
$s_{q}(x_{1},\dots,x_{n_{0}}):=\sum_{p=1}^{n_{0}}\psi_{pq}(x_{p}),\qquad q=1,\dots,2M+1.$ (3.3)
Theorem 3.2 states that every scalar-valued continuous function on $X$ is a finite sum of univariate superpositions $h_{q}\circ s_{q}$. In particular, if the functions $s_{q}$ are available as admissible features, then they yield a finite-dimensional composition in the sense of Definition 3.1. Indeed, letting
$F=(s_{1},\dots,s_{2M+1}):X\to\mathbb{R}^{2M+1},$
the exact superposition formula implies that every $g\in C(K;\mathbb{R}^{m})$ admits an exact representation of the form $g=u\circ F|_{K}$ for some $u\in C(\mathbb{R}^{2M+1};\mathbb{R}^{m})$. This is precisely the strong finite-dimensional composition property identified in Proposition 3.1.
Theorem 3.3 (Deep narrow universality from Ostrand-type features).
Let $X=\prod_{p=1}^{n_{0}}X_{p}$ be a product of compact metric spaces, with $\dim_{\mathrm{top}}(X_{p})=d_{p}<\infty$, and set $M=\sum_{p=1}^{n_{0}}d_{p}$. Let $\mathcal{A}(X)\subset C(X)$ be a feature family and assume that the Ostrand inner functions $s_{q}$ defined in (3.3) belong to $\mathcal{A}(X)$ for $q=1,\dots,2M+1$. Let $F=(s_{1},\dots,s_{2M+1})$ as above. Assume that $\sigma:\mathbb{R}\to\mathbb{R}$ is continuous and nonaffine, and that there exists $t_{0}\in\mathbb{R}$ and an open neighborhood $U$ of $t_{0}$ such that $\sigma$ is differentiable on $U$ and $\sigma^{\prime}$ is continuous and nonzero at $t_{0}$.
Then for every integer $m\geq 1$, every compact set $K\subset X$, every function $g\in C(K;\mathbb{R}^{m})$, and every $\varepsilon>0$, there exists a function
$H\in\mathcal{N}^{(2M+m+3,m)}(X,\sigma;F)$
such that
$\|g-H\|_{K}<\varepsilon.$
Equivalently, for each fixed $m$ and compact $K\subset X$, the class $\mathcal{N}^{(2M+m+3,m)}(X,\sigma;F)|_{K}$ is dense in $C(K;\mathbb{R}^{m})$.
Proof.
Fix $m\geq 1$, a compact set $K\subset X$, a target function $g=(g_{1},\dots,g_{m})\in C(K;\mathbb{R}^{m})$, and $\varepsilon>0$.
By Theorem 3.2, for each component $g_{r}\in C(K)$ ($r=1,\dots,m$) there exist functions $h_{r,q}\in C(\mathbb{R})$, $q=1,\dots,2M+1$, such that for all $x\in K$,
$g_{r}(x)=\sum_{q=1}^{2M+1}h_{r,q}\bigl(s_{q}(x)\bigr).$
Define $u:\mathbb{R}^{2M+1}\to\mathbb{R}^{m}$ by
$u(y_{1},\dots,y_{2M+1}):=\left(\sum_{q=1}^{2M+1}h_{1,q}(y_{q}),\ \dots,\ \sum_{q=1}^{2M+1}h_{m,q}(y_{q})\right).$
Then $u\in C(\mathbb{R}^{2M+1};\mathbb{R}^{m})$ and, by construction,
$g(x)=u(F(x))\qquad\text{for all }x\in K.$
Thus, $g$ admits an exact representation of the form $g=u\circ F|_{K}$, which is the strong finite-dimensional composition property from Proposition 3.1(i).
Since $F$ is continuous and $K$ is compact, the image $K_{F}:=F(K)$ is a compact subset of $\mathbb{R}^{2M+1}$. By the deep narrow universality theorem of Kidger and Lyons [18, Theorem 3.2], applied to the compact set $K_{F}$ and to the continuous map $u|_{K_{F}}\in C(K_{F};\mathbb{R}^{m})$, there exists a (classical) deep feedforward network
$\Phi:\mathbb{R}^{2M+1}\to\mathbb{R}^{m}$
with activation $\sigma$ and width at most $(2M+1)+m+2$ such that
$\max_{y\in K_{F}}\|u(y)-\Phi(y)\|_{\mathbb{R}^{m}}<\varepsilon.$
Define $H=\Phi\circ F$. Then, by Definition 3.2,
$H\in\mathcal{N}^{(2M+m+3,m)}(X,\sigma;F).$
Moreover, for $x\in K$ we have $F(x)\in K_{F}$, hence
$\|g(x)-H(x)\|_{\mathbb{R}^{m}}=\|u(F(x))-\Phi(F(x))\|_{\mathbb{R}^{m}}<\varepsilon.$
Taking the maximum over $x\in K$ yields $\|g-H\|_{K}<\varepsilon$. ∎
Remark 3.4.
In [14] we introduced topological feedforward neural networks and proved a universal approximation theorem for single-hidden-layer networks with scalar outputs on general topological spaces. As an application, we derived an approximative version of the Kolmogorov superposition theorem for compact metric spaces by replacing the non-fixed outer functions with a single infinitely smooth activation function.
The present paper extends this viewpoint in two directions. First, we treat vector-valued outputs and formulate universality within a unified framework encompassing both shallow and deep architectures. Second, we address the width-constrained setting: deep narrow universality is established under a finite-dimensional composition property, which identifies when approximation on a general topological space can be reduced to approximation on a Euclidean space. In particular, for products of compact metric spaces we obtain explicit width bounds expressed in terms of topological dimension, thereby linking geometric and topological structure to quantitative architectural constraints.
Remark 3.5.
Related results on approximation of mappings between general topological spaces are developed in [20]. In that work, one starts from a neural network class $\mathcal{F}\subset C(\mathbb{R}^{n},\mathbb{R}^{m})$ that is already known to be dense and investigates when density is preserved under modifications of the input and output layers. More precisely, under suitable assumptions on continuous maps $\phi:X\to\mathbb{R}^{n}$ and $\rho:\mathbb{R}^{m}\to Y$, in particular assuming that $\phi$ is injective, the authors obtain density of the composed class
$\{\,\rho\circ f\circ\phi:f\in\mathcal{F}\,\}$
in $C(X,Y)$ with respect to the topology of uniform convergence on compact sets.
Our deep narrow construction is related in spirit to composition-based transfer principles when $Y=\mathbb{R}^{m}$ and $\rho=\mathrm{id}$, since networks on $X$ are represented as compositions $H=\Phi\circ F$, where $\Phi$ is a Euclidean neural network and $F:X\to\mathbb{R}^{n}$ is a feature map. However, the role played by $F$ in the present work differs fundamentally from the role of the fixed inner map $\phi$ in [20]. In [20], the map $\phi$ is assumed a priori and serves as a fixed, globally injective mapping of $X$ into a Euclidean space, through which approximation properties are transferred from $\mathbb{R}^{n}$ to $X$.
In contrast, in our framework the feature map $F$ is intrinsic to the neural network architecture on $X$ and is constructed from an admissible feature family $\mathcal{A}(X)\subset C(X)$. In the universality results without width constraints, the maps $F=(f_{1},\dots,f_{n})$ entering the first hidden layer of the network representation depend on the compact set under consideration, the target function, and the required accuracy, similarly to trainable parameters in classical networks. In the deep narrow setting, a finite-dimensional composition is imposed, and $F=(f_{1},\dots,f_{n})$ depends only on the compact set $K$ and the output dimension $m$. Thus, in both the unconstrained and deep narrow settings, the feature map $F$ is variable rather than a single fixed global injective map as in [20].
Accordingly, in the present work injectivity of the feature map $F$ is not imposed a priori, but is shown to be a necessary consequence of the finite-dimensional composition property (Definition 3.1). This property specifies when a finite collection of admissible features $f_{1},\dots,f_{n}\in\mathcal{A}(X)$ suffices to reduce approximation on $K$ to approximation on the compact subset $F(K)\subset\mathbb{R}^{n}$. When $F|_{K}$ is a topological embedding, Proposition 3.1 shows that this reduction holds in a strong form, yielding exact representations $g=u\circ F|_{K}$ for all $g\in C(K;\mathbb{R}^{m})$.
This shifts the focus from the existence of an abstract embedding $\phi$ to concrete structural properties of the feature family $\mathcal{A}(X)$ and the topology of $X$. In particular, topological assumptions such as the Hausdorff property play a natural role: if $X$ is not Hausdorff, a feature map $F:X\to\mathbb{R}^{n}$ need not be a topological embedding, and the strong composition property may fail. Hence, our framework is inherently topology-sensitive rather than merely compositional.
Moreover, in the compact metric setting we give an explicit realizability mechanism based on classical dimension theory. Ostrand’s extension of the Kolmogorov superposition theorem yields concrete feature maps for products of compact metric spaces. This leads to deep narrow universality with width bounds stated in terms of topological dimension. In this way, geometric and topological structure translates into width-controlled universality on non-Euclidean spaces.
4 Discussion and Outlook
This work develops a general approximation-theoretic framework for neural networks whose inputs range over arbitrary topological spaces. By formulating neural architectures in terms of admissible feature families, we obtain a unified setting that encompasses classical feedforward networks on Euclidean spaces as a special case.
A central focus of the paper is the universal approximation property for both shallow and deep networks with vector-valued outputs. In the absence of width constraints, we establish general density results that extend classical universality theorems to non-Euclidean spaces.
The novelty also arises in the analysis of deep narrow networks, where the width of each hidden layer is uniformly bounded while depth is allowed to grow. We identify conditions under which such width-constrained networks retain universal approximation power, thereby extending deep narrow universality beyond the Euclidean setting.
An important outcome of this analysis is the explicit connection between universality of deep narrow networks and classical results from topology and dimension theory. In particular, Ostrand’s extension of the Kolmogorov superposition theorem yields concrete feature maps for products of compact metric spaces. As a consequence, we obtain explicit width bounds expressed in terms of topological dimension. This demonstrates how geometric and topological structure of the input space can be translated into approximation properties of the considered networks.
Several directions for further research remain open. One natural extension is the study of neural networks with outputs in infinite-dimensional spaces. Another promising direction is the investigation of quantitative approximation rates and complexity bounds within the present topological framework. Finally, it would be of interest to explore how the abstract feature-based approach developed here interfaces with practical applications.